OpenStudy (anonymous):

Describe the end behavior of polynomial graphs with odd and even degrees. Talk about positive and negative leading coefficients.

OpenStudy (anonymous):

think of the "end behaviour" of a line with positive or negative slope

OpenStudy (anonymous):

I just started this and my teacher told me to try this tonight and I don't understand it at all

OpenStudy (anonymous):

if the slope of a line is positive then as \(x\to \infty\) you have \(y\to \infty\) also

OpenStudy (anonymous):

and as \(x\to -\infty\) also \(y\to -\infty\) this is true of any odd polynomial with positive leading coefficient

OpenStudy (anonymous):

still very confused

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Those are the notes my teacher gave me.

OpenStudy (anonymous):

If a number has an even exponent, both sides of the graph are going to have the same end behaviour: either they will both go towards \[\infty\] or -\[\infty\]

OpenStudy (anonymous):

\[-\infty \]

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

If the number has an even exponent and a positive coefficient, that means both sides of the graph are going to head towards positive infinity. x→-∞ then y→∞ and x→∞ then y→∞

OpenStudy (anonymous):

Just for positive even exponent haha

OpenStudy (anonymous):

if the number has a negative coefficient and an even exponent then instead it will be x→-∞ then y→-∞ and x→∞ then y→-∞

OpenStudy (anonymous):

Like in those two graphs ^^

OpenStudy (anonymous):

Now if a number has an odd exponent, one side of the graph will go towards negative infinity and one side will go towards positive infinity

OpenStudy (anonymous):

If its a negative coefficient and odd exponent, it'll look like this:

OpenStudy (anonymous):

so first it heads towards positive infinity and then negative infinity. x→-∞ then y→∞ and x→∞ then y→-∞

OpenStudy (anonymous):

but if its an odd exponent and positive coefficient, it would be x→-∞ then y→-∞ and x→∞ then y→∞

OpenStudy (anonymous):

Sooo 2x^4 x→-∞ then y→∞ and x→∞ then y→∞ -2x^4 x→-∞ then y→-∞ and x→∞ then y→-∞ 2x^3 x→-∞ then y→-∞ and x→∞ then y→∞ -2x^3 x→-∞ then y→∞ and x→∞ then y→-∞

OpenStudy (anonymous):

@thenikko so it would be positive goes up and negative goes down?

OpenStudy (anonymous):

For even exponents

OpenStudy (anonymous):

@thenikko so what would the answer be? For positive exponents positive goes up and negative goes down. But with an odd exponent the both go towards positive?

OpenStudy (anonymous):

For positive, even exponents both go up. For negative, even exponents both go down. For positive, odd exponents it goes down and then up. For negative, odd exponents it goes up and then down